Problem 7-05 Nonconstant Growth Valuation A company currently pays a dividend of $2.5 per share (D0 = $2.5). It is estimated that the company's dividend will grow at a rate of 15% per year for the next 2 years, and then at a constant rate of 8% thereafter. The company's stock has a beta of 1.4, the risk-free rate is 5.5%, and the market risk premium is 5%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent. $
Answer #1

HI

At first we will calculate required rate of return for the company using CAPM

required return k = rf+beta*(rm-rf)

= 5.5 + 1.4*5

= 12.5%

Now First year Dividend D1 = 2.5*(1+15%)= $2.875

Second year dividend =2.875*(1+15%) = $3.30625

Third year dividend D3 = 3.30625*(1+8%) = $3.57075

Now Present value of dividends = D1/(1+k) +D2/(1+k)^2 + (D3/(k-g))/(1+k)^2

=2.875/(1+12.5%) + 3.30625/(1+12.5%)^2 + 3.57075/(12.5%-8%)/(1+12.5%)^2

= $67.86

Thanks